Semi-annual payment of 12,000.00 for twelve (12) years that will start from now. What is the period of deferral?


1
Expert's answer
2022-01-18T11:13:11-0500

Solution:

"\begin{aligned}

&\text { Semi-Annual Payments }=\mathrm{P} 12,000 \\

&\text { Time Period for which Payment is made }=12 \text { Years } \\

&\text { Payment start made after } n\text {(missing in question)} \text { Years from now } \\

&\text { Number of semi-annual periods in a year }=2 \\

&\text { Number of deferral periods }=\text { Number of Semi Annual Periods } \times \text { Time Period after which payment starts(t) - } 1 \\

&\Rightarrow\text { Number of deferral periods }=2 \times n-1 \\

&\Rightarrow\text { Number of deferral periods }=2n-1 \\

\end{aligned}"